OpenStudy (anonymous):

Find all rational zeros of the polynomial P(x)=x^4-3x^3-2x^2-6x-8 Enter the rational zeros in a comma separated list.

OpenStudy (anonymous):

Are you sure this is the exact question and all the signs are correct?

OpenStudy (anonymous):

Unfortunately Yes

OpenStudy (anonymous):

I think I got it: f(x)=0= x^4-3x^3-2x^2-6x-8 Factor by grouping: (x^4-2x^2-8)-(3x^3-6x) = 0

OpenStudy (anonymous):

Missed a sign: x^4-2x^2-8)-(3x^3+6x) = 0

OpenStudy (anonymous):

Try to factor this and tell me what you get

OpenStudy (anonymous):

Are you there?

OpenStudy (anonymous):

Im here @philo1234

OpenStudy (anonymous):

Ok try to factor the last expression by grouping that I gave. and tell me what you got

OpenStudy (anonymous):

(x^4-2x^2-8)-(3x^3+6x) = 0

OpenStudy (anonymous):

(x+1)(x-4)(x^2+2)

OpenStudy (anonymous):

There it is. good job.

OpenStudy (anonymous):

so 1, 4 and sqrt 2 is the answers

OpenStudy (anonymous):

\[-1,4, \pm i \sqrt{2}\]